OpenStudy (avai70178):

Please simplify. (4a^2b)^3(3ab^3)^2

OpenStudy (avai70178):

The little arrows pointing upwards mean the number to its right is a exponent

OpenStudy (anonymous):

\[576 a^8 b^9\]

OpenStudy (avai70178):

how did you get that

OpenStudy (avai70178):

i got 64a^5b^4

OpenStudy (anonymous):

\[(4 a^2 b^)3 (3 a b^3)^2\]

OpenStudy (anonymous):

Umm then dont trust me :D if you think you got it right

OpenStudy (avai70178):

lol ok thank you though

OpenStudy (avai70178):

do you want the medal?

OpenStudy (***[isuru]***):

lol.... blitzkrig is correct

OpenStudy (avai70178):

how

OpenStudy (anonymous):

MWHAHA I AM AWESOME ty ISURU

OpenStudy (anonymous):

eh? whos with me

OpenStudy (***[isuru]***):

hmm............ expand \[(4a^{2}b)^{3}\] u will get \[4^{3} a^{2 \times 3} b^{3}\] which is \[64a^{6}b^{3}\]

OpenStudy (anonymous):

thank you sasuke i honor you :)

OpenStudy (anonymous):

does that mean i got it wrong :OO man i suck *hides in corner*

OpenStudy (avai70178):

lol we were both wrong ty!

OpenStudy (***[isuru]***):

now expand \[(3ab^{3})^{2}\] which will lead u to \[3^{2} a^{2}b^{3 \times 2}\] which is \[9a^{2}b^{6}\]

OpenStudy (anonymous):

Sasuke you have beat me *slides away*

OpenStudy (***[isuru]***):

now multiply \[9a^{2}b^{6} \times 64a^{6}b^{3}\]

OpenStudy (anonymous):

so am i right?

OpenStudy (***[isuru]***):

the powers of like terms get add together

OpenStudy (***[isuru]***):

9 x 64 = 576

OpenStudy (***[isuru]***):

so... the answer is........ \[576a^{(2 + 6)}b^{(6+3)}\]

OpenStudy (***[isuru]***):

\[576a^{8}b^{9}\]

OpenStudy (anonymous):

YES IM AWESOME

OpenStudy (avai70178):

Thanks guys!

OpenStudy (***[isuru]***):

u r welcome!